Where do I begin with this equation?
The value of a home bought in 1990 appreciates, or increases as time passes. Seven years after it was bought, the house was worth $165,000. Twelve years after the purchase of the house, the house was worth $180,000.

a. If the relationshi is LINEAR, write an equation that describes this relationship.(HINT: Start by writing two ordered pairs!)

b. Estimate the value of the house in the year 2005, which is 15 years after the purchase of the home.

Answer by stanbon(75887)

The value of a home bought in 1990 appreciates, or increases as time passes. Seven years after it was bought, the house was worth $165,000. Twelve years after the purchase of the house, the house was worth $180,000.
-----------------------
You have two points (7,165000), (12,180000)
slope = 15000/5 = 3000
---------------
165000 = 3000*7 + b
b = 144000
-----------
EQUATION:
value = 3000(years after purchase) + 144000
---------------
b. Estimate the value of the house in the year 2005, which is 15 years after the purchase of the home.
V(15) = 3000*15 + 144000
V(15) = 45000+144000
V(15) = 189000
=====================
Cheers,
San H.